OpenStudy (anonymous):

Maria looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 13 inches. The length of the diagonal of the floor of the room in the plan is approximately 15.26 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

OpenStudy (emogirlatemoooocow):

is there answer choices ? \m/(>.<)\m/

OpenStudy (anonymous):

No

OpenStudy (emogirlatemoooocow):

okay just making sure

OpenStudy (tacotime):

Use the Pythagorean theorem. c^2=a^2+b^2 The hypotenuse is 15.26 and one side is 15. If the remaining side is 15 then it is a square.

OpenStudy (emogirlatemoooocow):



OpenStudy (anonymous):

ok

OpenStudy (emogirlatemoooocow):

I think you need to solve this with the Pythagorean theorem

OpenStudy (emogirlatemoooocow):

do you know how to sovle this with Pythagorean theorem ?

OpenStudy (emogirlatemoooocow):

@thuglife187

OpenStudy (anonymous):

ummm maybe?

OpenStudy (anonymous):

if the floor is a square, then the other side will also be 13 right ?

OpenStudy (anonymous):

Yes because all sides of a square are equal.

OpenStudy (anonymous):

@EmogirlAtEmoooocow is right with the picture

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

 but she need to add the 8 like this V

OpenStudy (emogirlatemoooocow):

oh lol but It says It's approximately 15.26 inches. @thuglife187

OpenStudy (anonymous):

\[13^{2} + 13^{2}\] isn't 15.26 inches then it isn't a square

OpenStudy (anonymous):

yes, however if we assume the floor as a square you get diagonal

OpenStudy (anonymous):

@EmogirlAtEmoooocow i send you the picture can you Attach it ?

OpenStudy (emogirlatemoooocow):

okay

OpenStudy (anonymous):

so 13^2 + 13^2 =338 and the square root of 338 is 18.38477631 so it is not a square

OpenStudy (anonymous):

yes